OpenStudy (anonymous):

Inverse of 1+x/(2+x)?

OpenStudy (anonymous):

y = 1 + x/(2+x) y(2+x) = 2 + x + x 2y + xy = 2 + 2x xy - 2x = 2 - 2y x (y-2) = 2 - 2y x = 2(1-y) / (y-2) inverse = f-1(x) = 2(1-x) / (x-2)

OpenStudy (anonymous):

the answer sheet says the denominantor is 1-x

OpenStudy (anonymous):

hmm - cant see whats wrong with my algebra

OpenStudy (anonymous):

x=1+y/2+y x(2+y)=1+y 2x+xy=1+y xy-y=-2x+1 y(x-1)=-2x+1 y=-2x+1/x-1

OpenStudy (anonymous):

you had one to many x's in your second step

OpenStudy (anonymous):

aaaand the numerator isnt -2x on the answer sheet

OpenStudy (anonymous):

hmmm, now Im a little confused. lol

OpenStudy (anonymous):

oh i see whats happened - i assumed that original expression was 1 + (x / 2+x) - it would avoid confusion if you had typed it as (1+x)/ (2 + x)

OpenStudy (anonymous):

the answer is 2x-1/1-x...

OpenStudy (anonymous):

oooh sorry

OpenStudy (anonymous):

Ahh ok so they factored out a negative so it should look like this 2x-1/-x+1

OpenStudy (anonymous):

or the way you wrote they are equal

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

gotchya